# leftmargin=1em, rightmargin=1em, indentfirst=false

### v2

@@ Abstract
Large language models (LLMs) have emerged as a dominant AI paradigm due to their exceptional text understanding and generation capabilities. However, their tendency to generate inconsistent or erroneous outputs challenges their reliability, especially in high-stakes domains requiring accuracy and trustworthiness. Existing research primarily focuses on detecting and mitigating model misbehavior in general-purpose scenarios, often overlooking the potential of integrating domain-specific knowledge. In this work, we advance misbehavior detection by incorporating domain knowledge. The core idea is to design a general specification language that enables domain experts to customize domain-specific predicates in a lightweight and intuitive manner, supporting later runtime verification of LLM outputs. To achieve this, we design a novel specification language ESL and introduce a runtime verification framework RvLLM to validate LLM output against domain-specific constraints defined in ESL. RvLLM operates in two main stages: interpretation and reasoning. During interpretation, it derives interpretations of the specification based on context, which then guide the reasoning process to identify inconsistencies. When new knowledge is derived, RvLLM issues a follow-up query to the LLM to further verify the consistency. We evaluate RvLLM on three representative tasks: violation detection against Singapore Rapid Transit Systems Act, numerical comparison, and inequality solving. Experimental results demonstrate that RvLLM effectively detects erroneous outputs across various LLMs in a lightweight and flexible manner. The results reveal that despite their impressive capabilities, LLMs remain prone to low-level errors due to limited interpretability and a lack of formal guarantees during inference, and our framework offers a potential long-term solution by leveraging expert domain knowledge to rigorously and efficiently verify LLM outputs.
@@ 1 Introduction
Unlike rule-based systems [47] operating on predefined and deterministic rules, large language models (LLMs) [3, 21, 56, 39] learn data representation and processing automatically from the training datasets, achieving human-like or even superhuman performance, and have driven significant advancements in various practical applications [11, 30, 50, 40, 22]. However, their non-deterministic and unpredictable nature sometimes leads to inconsistent or erroneous outputs [5, 45], posing significant risks in safety-critical or knowledge-intensive domains. Recent studies [55, 6] have shown that such misbehavior in LLMs cannot be fully eliminated, underscoring the urgent need for dedicated verification and validation methodologies to enhance the reliability of LLM-generated outputs.
@@
LLM testing [59, 25, 28, 60] primarily aims to establish comprehensive benchmarks to evaluate the overall model performance against domain-agnostic criteria–such as accuracy, coherence, and fairness–in alignment with the intended application. While these approaches effectively assess general behavior and reveal edge cases that may provoke unexpected responses, they are limited to predefined benchmarks and lack the specificity needed to address domain-specific assessment needs. LLM verification, instead, may serve as a complementary mechanism to LLM testing by providing formal guarantees on model behavior. Despite substantial progress over the past decade in neural network verification [9, 17, 27, 31, 54, 44, 57], existing methods exhibit notable limitations: they are primarily tailored to simpler model architectures–such as deep or convolutional networks–and classification tasks, struggling to scale to the complexity of modern AI models and their diverse functionalities. Furthermore, these approaches target at general properties such as robustness and fairness, making them ill-suited for domain-specific properties. Therefore, developing a dedicated verification paradigm tailored to LLMs is crucial for ensuring certified and reliable outputs.
@@
On the other hand, the decline of rule-based expert systems [23, 29, 20] in the late 20th century can be attributed to their inability to handle incomplete information and poor scalability in real-world complexity. Defining comprehensive rule sets for open domains has been proven infeasible [47, 26]. We argue that such rules can only be statistically approximated–a capability exemplified by LLMs. However, this approximation cannot substitute for explicit rule encoding, and learning-based models remain inherently unreliable, particularly in edge cases demanding strict adherence to domain-specific requirements. This highlights the need for verifying explicitly defined rules in a lightweight and incremental manner for practical deployment and maintainability.
Figure 1: An overview of RvLLM. Given the LLM’s context and outputs, RvLLM first extracts relevant propositions and translates the rules from the specification file into a normalized set of propositional logic formulae. These formulae are then processed via forward chaining to detect inconsistencies or infer new knowledge, which is then used to generate follow-up queries and assess consistency across successive LLM responses.
@@
This paper. Building on these foundations, we propose runtime verification with domain knowledge as a sustainable solution to ensuring reliable LLM behavior. Our work is motivated by two key insights: i) Runtime verification offers a lightweight yet rigorous means to bridge testing and formal verification by assessing system behavior against formal properties during execution; ii) Existing work primarily focuses on general misbehavior detection, and often overlooks the critical role of domain-specific expertise. Integrating such knowledge is essential for handling edge cases and improving LLM reliability in specialized tasks, as domain knowledge can significantly enhance model performance.
@@
To achieve this, we introduce RvLLM, an innovative runtime verification framework tailored to LLMs to ensure reliable outputs, particularly against axiomatic domain specifications. The idea is to automatically check whether the LLM-generated responses adhere to domain-specific properties expressed in a simple, adaptable specification language. This simplicity and flexibility empower domain experts to encode their knowledge and formally define constraints that capture the expected behavior of LLMs in specialized applications. To support this, we introduce ESL, a general language tailored for encoding rule-based domain expertise. ESL integrates natural language with formal logic to ensure the adaptability to the natural language settings of LLMs while enabling the structured and formal representation of properties to verify. This design allows for the rigorous specification of domain-specific criteria across diverse LLM applications.
@@
Experimental results. We evaluate the effectiveness of RvLLM through experiments on three representative tasks: violation detection against Singapore Rapid Transit Systems Act [1], numerical comparison, and inequality solving, using a diverse set of LLMs. The experimental results show that RvLLM significantly enhances the reliability of LLM output in these domain-specific tasks. In the violation detection task, employing RvLLM as a complementary mechanism increases the true positive rate, or TPR, by 15.7% to 50.2% across various models. In the numerical comparison task, RvLLM detects nearly all errors in the LLM responses across 100 randomly generated questions. In the inequality solving task, RvLLM facilitates a systematic analysis of the step-by-step reasoning process from LLMs, effectively identifying its deviations from expert-defined specifications. These specifications are primarily grounded in fundamental algebraic inequality properties typically covered at the junior college level. Notably, despite the incomplete domain knowledge in the inequality solving task, RvLLM still achieves a TPR of 52.2% in detecting erroneous solutions. We argue that such a runtime verification approach can function as a critical safeguard for LLMs, particularly in rule-based or reasoning-intensive domains where adherence to specific constraints is essential.
2 Related work
LLM testing. Research in this area aims to develop comprehensive benchmarks and systematic evaluation frameworks to assess LLM across general attributes such as robustness, fairness, and factual consistency [41, 37, 52]. Robustness studies examine model resilience to adversarial input, paraphrasing, or noise [10, 51, 58], while fairness evaluations assess bias with respect to sensitive attributes like gender or race [35, 58]. Factual consistency detection, or hallucination detection, focuses on identifying instances in which LLMs produce plausible yet factually incorrect or unsupported claims [38, 34, 36]. Although large-scale benchmarks like HELM [37] and BIG-bench [49] offer task suites and metrics for rigorous performance evaluation, spanning accuracy, calibration, and ethical considerations, a critical challenge persists, i.e., the dynamic and open-ended nature of real-world deployments complicates the creation of exhaustive and representative test suites.
Runtime verification of LLMs. Runtime verification [7, 19] is a dynamic analysis technique that ensures certified system behavior during execution by monitoring traces against formal specifications. Owing to its lightweight nature, runtime verification has inspired multiple efforts to verify LLM-generated outputs at runtime [14, 8, 43, 12, 33, 42, 32]. However, existing approaches center on open-domain settings, where specifications are often ambiguous and informal. For instance, SelfCheckGPT [43] and CHECKEMBED [8] operationalize specifications in terms of output stability. More recently, [13] proposes a fairness monitoring approach leveraging precisely defined properties in formal specifications. These specifications, informed by linear temporal logic [46] and its bounded metric variant [4], signal a shift towards formal methods for more accurate and dependable runtime verification of LLM behaviors. However, these works primarily focus on general properties and are not well suited to domain-specific constraints for specialized tasks.
@@ 3.1 Design guidance
In the following, we give our high-level requirements for designing a specification language that encapsulates LLM behavioral constraints from the perspective of domain experts. The core objective is to strike a balance between user-friendliness and expressiveness. We illustrate this with the following regulation from Singapore Rapid Transit System Act [1]: {quoting} “No person shall consume or attempt to consume any chewing gum or bubble gum while in or upon any part of the railway premises.”
Accessibility for domain experts. The primary goal of this specification language is to empower domain experts, even those with limited programming or formal methods experience, to effectively define and enforce constraints on the LLM’s behavior. To balance expressiveness and user-friendliness, we base the specification rule on predicate logic [48, 18], restricted to prenex normal forms [48, 18] and consisting solely of universal quantifiers. We further constrain its quantifier-free part to a rule-like form (or implication), yielding a simplified variant of predicate logic expressible as $\forall\vec{x}.(f(\vec{x})\Rightarrow g(\vec{x}))$ [^1] We use $\vec{x}$ instead of $x$ as there could be a sequence of variables.. Specifically, both the left-hand side (LHS) and right-hand side (RHS) of the rule-like part are quantifier-free predicate formulas composed of predicates applied to terms that may include variables (e.g., $x$), constants (e.g., ‘bubble gum’), and functions of terms (e.g., $\text{sin}(x)$). For example, ${\sf ChewGum}(x)$ is an atomic predicate with a single variable argument $x$, and ${\sf IsGreater}({\sf Square}(x),0)$ is also an atomic predicate with two arguments: the function term ${\sf Square}(x)$ and the constant term 0.
@@ 3.2 Formalization
Given a propositional formula $\psi_{1}\Rightarrow\psi_{2}$, if $\psi_{1}$ is a disjunctive norm form [48] and $\psi_{2}$ is a conjunctive norm form [48], then we call $\psi_{1}\Rightarrow\psi_{2}$ is a deductive normal form (DeNF).
@@
An ESL rule is a DeNF where each variable is substituted by a predicate.
@@
{minted}
[fontsize=,bgcolor=backgroundcolor,escapeinside=||]python "Variables": "x", "Predicates": "ChewGum(x) := a person x consumes or attempts to consume chewing gum or bublle gum.", "InRailway(x) := a person x is in some part of railway premises", "Rules": "InRailway(x) => not ChewGum(x)"
[…]
@@ 4.1 Interpretation abstraction
Consider the example in Figure 2, which illustrates our approach using a simple numerical comparison question with a specification encoding a basic algebraic property of multiplication, the interpretation abstraction process first gets all possible objects and propositionalizes the predicate ${\sf IsGreater}$ as follows: $\{p_{0}={\sf IsGreater}(o_{1},o_{2})={\sf True},p_{1}={\sf IsGreater}(o_{2},o_{1})={\sf False}\}$ with $o_{1}=15.2$ and $o_{2}=15.12$. Then, it can obtain all possible variable binding $\{\flat_{1},\flat_{2}\}$ for the rule $\psi$ where $\flat_{1}=\{x\mapsto o_{1},y\mapsto o_{2}\}$ and $\flat_{2}\{x\mapsto o_{2},y\mapsto o_{1}\}$. $\flat_{1}$ results in a formula $p_{0}\wedge{\sf IsGreater}(z,0)\Rightarrow{\sf IsGreater}(15.2*z)$ for $\psi$, which is an partial binding. Similarly, $\flat_{2}$ is also partial. Consequently, for the Level-1 interpretation, no propositional formula is successfully generated, and the checking procedure exits with no inconsistency detected. While for the Level-2 interpretation, it successfully instantiates the variable $z$ for the partial biddings $\flat_{1}$ and $\flat_{2}$ with an object $o_{3}=10$, and obtain an additional proposition $p_{2}={\sf IsGreater}(o_{3},0)={\sf True}$. Then, the updated bindings are: $\flat_{1}^{\prime}=\{x\mapsto o_{1},y\mapsto o_{2},z\mapsto o_{3}\},\flat_{2}^{\prime}=\{x\mapsto o_{2},y\mapsto o_{1},z\mapsto o_{3}\}$. By applying $\flat_{1}^{\prime}$, we can obtain a new proposition $p_{3}={\sf IsGreater}(o_{1}*o_{3},o_{2}*o_{3})={\sf IsGreater}(15.2*10,15.12*10)$ but with unknown truth value. Similarly, we obtain $p_{4}={\sf IsGreater}(o_{2}*o_{3},o_{1}*o_{3})={\sf IsGreater}(15.12*10,15.2*10)={\sf Unknown}$ by $\flat_{2}^{\prime}$. Finally, we generate two propositional formulae: $\psi_{1}=p_{0}\wedge p_{2}\Rightarrow p_{3}$ via $\flat_{1}^{\prime}$ and $\psi_{2}=p_{1}\wedge p_{2}\Rightarrow p_{4}$ via $\flat_{2}^{\prime}$ for the following component.

### v3

# RvLLM: LLM Runtime Verification with Domain Knowledge

@@ Abstract
Large language models (LLMs) have emerged as a dominant AI paradigm due to their exceptional text understanding and generation capabilities. However, their tendency to generate inconsistent or erroneous outputs challenges their reliability, especially in high-stakes domains requiring accuracy and trustworthiness. Existing research primarily focuses on detecting and mitigating model misbehavior in general-purpose scenarios, often overlooking the potential of integrating domain-specific knowledge. In this work, we advance misbehavior detection by incorporating domain knowledge. The core idea is to design a general specification language that enables domain experts to customize domain-specific constraints in a lightweight and intuitive manner, supporting later runtime monitoring of LLM outputs. To achieve this, we design a novel specification language ESL and introduce a runtime verification framework RvLLM to validate LLM output against domain-specific constraints defined in ESL. RvLLM operates in two main stages: interpretation and reasoning. During interpretation, it derives interpretations of the specification based on the context, which then guide the reasoning process to identify inconsistencies. When new knowledge is derived, RvLLM issues a follow-up query to the LLM to further verify the consistency. We evaluate RvLLM on three representative tasks: violation detection against Singapore Rapid Transit Systems Act, numerical comparison, and inequality solving. Experimental results show that RvLLM effectively detects erroneous outputs across various LLMs in a lightweight and flexible manner. The results reveal that despite their impressive capabilities, LLMs remain prone to low-level errors due to a lack of formal guarantees during inference, and our framework offers a potential long-term solution by leveraging expert domain knowledge to rigorously and efficiently verify LLM outputs.
@@ 1 Introduction
Unlike rule-based systems [russell2016artificial] operating on predefined and deterministic rules, large language models (LLMs) [achiam2023gpt, geminiteam2024gemini, yang2024qwen2, liu2024deepseek] learn data representation and processing automatically from the training datasets, achieving human-like or even superhuman performance, and have driven significant advancements in various practical applications [cascella2023evaluating, kasneci2023chatgpt, thirunavukarasu2023large, liu2023your, guha2023legalbench]. However, their non-deterministic and unpredictable nature sometimes leads to inconsistent or erroneous outputs [lawBBC, suicideBBC], posing significant risks in safety-critical or knowledge-intensive domains. Recent studies [xu2024hallucination, banerjee2024llms] have shown that such misbehavior in LLMs cannot be fully eliminated, underscoring the urgent need for dedicated verification and validation methodologies to enhance the reliability of LLM-generated outputs.
@@
LLM testing [zhong2024agieval, HendrycksBBZMSS21, huang2024c, zhou2023don] primarily aims to establish comprehensive benchmarks to evaluate the overall model performance against domain-agnostic criteria–such as accuracy, coherence, and fairness–in alignment with the intended application. While these approaches effectively assess general behavior and reveal edge cases that may provoke unexpected responses, they are limited to predefined benchmarks and lack the specificity needed to address domain-specific assessment needs. LLM verification, instead, may serve as a complementary mechanism to LLM testing by providing formal guarantees on model behavior. Despite substantial progress over the past decade in neural network verification [BunelLTTKK20, ElboherGK20, HKWW17, KBDJK17, betaCrown, narodytska2018verifying, BDD4BNN], existing methods exhibit notable limitations: they are primarily tailored to simpler model architectures–such as deep or convolutional networks–and classification tasks, struggling to scale to the complexity of modern AI models and their diverse functionalities. Furthermore, these approaches target general properties such as robustness and fairness, making them ill-suited for domain-specific properties. Therefore, developing a dedicated verification paradigm tailored to LLMs is crucial for ensuring certified and reliable outputs.
@@
On the other hand, the decline of rule-based expert systems [hayes1983building, jackson1986introduction, forsyth1984expert] in the late 20th century can be attributed to their inability to handle incomplete information and poor scalability in real-world complexity. Defining comprehensive rule sets for open domains has been proven infeasible [russell2016artificial, hernandez2017measure]. We argue that such rules can only be statistically approximated–a capability exemplified by LLMs. However, this approximation cannot substitute for explicit rule encoding, and learning-based models remain inherently unreliable, particularly in edge cases demanding strict adherence to domain-specific requirements. This highlights the need for verifying explicitly defined rules in a lightweight and incremental manner for practical deployment and maintainability.
@@
This paper. Building on these foundations, we propose runtime verification with domain knowledge as a sustainable solution to ensuring reliable LLM behavior, motivated by two key insights: i) Runtime verification offers a lightweight yet rigorous means to bridge testing and formal verification by assessing system behavior against formal properties during execution; ii) Existing work primarily focuses on generic misbehavior detection while often overlooking the critical role of domain-specific expertise. Integrating such knowledge is crucial for handling edge cases and enhancing LLM reliability in specialized tasks, where domain knowledge can significantly improve performance.
@@
To achieve this, we introduce RvLLM, an innovative runtime verification framework tailored to LLMs to ensure reliable outputs, particularly against axiomatic domain specifications. The idea is to automatically check whether the LLM-generated responses adhere to domain-specific constraints expressed in a simple, adaptable specification language. This simplicity and flexibility empower domain experts to encode their knowledge and formally define constraints that capture the expected behavior of LLMs in specialized applications. To support this, we introduce ESL, a general language tailored for encoding rule-based domain expertise. ESL integrates natural language with formal logic to ensure the adaptability to the natural language settings of LLMs while enabling the structured and formal representation of properties to verify. This design allows for the rigorous specification of domain-specific criteria across diverse LLM applications.
@@
Figure 1: An overview of RvLLM. Given an LLM’s context and outputs, RvLLM first extracts relevant propositions and translates the ESL specification into normalized propositional logic formulae. These are then processed through forward chaining to detect inconsistencies and infer new knowledge, which subsequently guides follow-up queries and consistency checks across successive LLM responses.
Experimental results. We evaluate the effectiveness of RvLLM through experiments on three representative tasks: violation detection against Singapore Rapid Transit Systems Act [SGMRT], numerical comparison, and inequality solving, using a diverse set of LLMs. The experimental results show that RvLLM significantly enhances the reliability of LLM output in these domain-specific tasks. In the violation detection task, employing RvLLM as a complementary mechanism increases the true positive rate (TPR) by 15.7% to 50.2% across various models. In the numerical comparison task, RvLLM detects nearly all errors in the LLM responses across 100 randomly generated questions. In the inequality solving task, RvLLM facilitates a systematic analysis of the step-by-step reasoning process from LLMs, effectively identifying its deviations from expert-defined specifications. These specifications are primarily grounded in fundamental algebraic inequality properties typically covered at the junior college level. Notably, despite the incomplete domain knowledge in the inequality solving task, RvLLM still achieves a TPR of 50% in detecting erroneous solutions. We argue that such a runtime verification approach can function as a critical safeguard for LLMs, particularly in rule-based or reasoning-intensive domains where adherence to specific constraints is essential. We further compare RvLLM with existing runtime hallucination-detection methods. Results show that RvLLM demonstrates superior performance by maintaining both high true positive and negative rates, whereas existing methods often exhibit a non-negligible trade-off between these metrics.
@@ 2.1 Design guidance
In the following, we give our high-level requirements for designing a specification language that encapsulates LLM behavioral constraints from the perspective of domain experts. The core objective is to strike a balance between user-friendliness and expressiveness. We illustrate this with the following regulation from Singapore Rapid Transit System Act [SGMRT]:
  • “No person shall consume or attempt to consume any chewing gum or bubble gum while in or upon any part of the railway premises.”
Accessibility for domain experts. The primary goal of this specification language is to empower domain experts, even those with limited programming or formal methods experience, to effectively define and enforce constraints on the LLM’s behavior. To balance expressiveness and user-friendliness, we base the specification rule on predicate logic [shoenfield2018mathematical, enderton2001mathematical], restricted to prenex normal forms [shoenfield2018mathematical, enderton2001mathematical] and consisting solely of universal quantifiers. We further constrain its quantifier-free part to a rule-like form (or implication), yielding a simplified variant of predicate logic expressible as $\forall\vec{x}.(f(\vec{x})\Rightarrow g(\vec{x}))$ [^1] We use $\vec{x}$ instead of $x$ as there could be a sequence of variables.. Specifically, both the left-hand side (LHS) and right-hand side (RHS) of the rule-like part are quantifier-free predicate formulae composed of predicates applied to terms that may include variables (e.g., $x$), constants (e.g., “bubble gum”), and functions of terms (e.g., $\text{sin}(x)$). For example, ${\sf ChewGum}(x)$ is an atomic predicate with a single variable argument $x$, and ${\sf IsGreater}({\sf Square}(x),0)$ is also an atomic predicate with two arguments: the function term ${\sf Square}(x)$ and the constant term 0.
@@ 2.2 Formalization
Given a propositional formula $\psi_{1}\Rightarrow\psi_{2}$, if $\psi_{1}$ is a disjunctive norm form [shoenfield2018mathematical] and $\psi_{2}$ is a conjunctive norm form [shoenfield2018mathematical], then we call $\psi_{1}\Rightarrow\psi_{2}$ is a deductive normal form (DeNF).
@@
An ESL rule is a DeNF where each proposition is substituted by a predicate.
@@
"Variables": {"x"},
"Predicates": {"ChewGum(x) := a person x consumes or attempts to consume chewing gum or bubble gum.",
"InRailway(x) := a person x is in some part of railway premises"},
"Rules": {"InRailway(x) => not ChewGum(x)"}
Interpretation of ESL rules. An interpretation of an ESL rule $r$ assigns objects to variables (i.e., variable binding) in the predicate of $r$ and evaluates the truth value ($\sf True$}, $\sf False$}, or $\sf Unknown$}) of the resulting proposition after substitution. Note that we adopt the open-world assumption [drummond2006open] to better accommodate the open-ended nature of LLMs. In this work, interpretations are derived from the context and LLM outputs, serving as the domain of discourse [hein2003discrete]. Now, consider the following contextual scenario:
  • “In a crowded MRT train, Alex nervously chews gum to ease stress before an interview. Suddenly, the train jolts, and the gum flies out, landing on a stranger’s shirt. Awkward glances turn into laughter as apologies spill out, diffusing tension in the confined space.”
The interpretation of the rule ${\sf InRailway}(x)\Rightarrow\neg{\sf ChewGum}(x)$ is $\{{\sf InRailway}(o_{1})={\sf True}$, ${\sf InRailway}(o_{2})={\sf True}$, ${\sf ChewGum}(o_{1})={\sf True}$, ${\sf ChewGum}(o_{2})={\sf Unknown}\}$, where $o_{1}=\text{`Alex'}$ and $o_{2}=\text{`stranger'}$.
@@ 3.1 Interpretation abstraction
Consider the example in Figure 2, which illustrates our approach using a simple numerical comparison question with a specification encoding a basic algebraic property of multiplication, the interpretation abstraction process first gets all possible objects and propositionalizes the predicate $\sf IsGreater$} as follows: $\{p_{0}={\sf IsGreater}(o_{1},o_{2})={\sf True},p_{1}={\sf IsGreater}(o_{2},o_{1})={\sf False}\}$ with $o_{1}=15.2$ and $o_{2}=15.12$. Then, it can obtain all possible variable binding $\{\flat_{1},\flat_{2}\}$ for the rule $\psi$ where $\flat_{1}=\{x\mapsto o_{1},y\mapsto o_{2}\}$ and $\flat_{2}\{x\mapsto o_{2},y\mapsto o_{1}\}$. $\flat_{1}$ results in a formula $p_{0}\wedge{\sf IsGreater}(z,0)\Rightarrow{\sf IsGreater}(15.2*z)$ for $\psi$, which is an partial binding. Similarly, $\flat_{2}$ is also partial. Consequently, for the Level-1 interpretation, no propositional formula is successfully generated, and the checking procedure exits with no inconsistency detected. While for the Level-2 interpretation, it successfully instantiates the variable $z$ for the partial bindings $\flat_{1}$ and $\flat_{2}$ with an object $o_{3}=10$, and obtain an additional proposition $p_{2}={\sf IsGreater}(o_{3},0)={\sf True}$. Then, the updated bindings are: $\flat_{1}^{\prime}=\{x\mapsto o_{1},y\mapsto o_{2},z\mapsto o_{3}\},\flat_{2}^{\prime}=\{x\mapsto o_{2},y\mapsto o_{1},z\mapsto o_{3}\}$. By applying $\flat_{1}^{\prime}$, we can obtain a new proposition $p_{3}={\sf IsGreater}(o_{1}*o_{3},o_{2}*o_{3})={\sf IsGreater}(15.2*10,15.12*10)$ but with unknown truth value. Similarly, we obtain $p_{4}={\sf IsGreater}(o_{2}*o_{3},o_{1}*o_{3})={\sf IsGreater}(15.12*10,15.2*10)={\sf Unknown}$ by $\flat_{2}^{\prime}$. Finally, we generate two propositional formulae: $\psi_{1}=p_{0}\wedge p_{2}\Rightarrow p_{3}$ via $\flat_{1}^{\prime}$ and $\psi_{2}=p_{1}\wedge p_{2}\Rightarrow p_{4}$ via $\flat_{2}^{\prime}$ for the following component.